### v5

# LangForce: Bayesian Decomposition of Vision Language Action Models via Latent Action Queries

Abstract
Vision-Language-Action (VLA) models have shown promise in robot manipulation but often struggle to generalize to new instructions or complex multi-task scenarios. We identify a critical pathology in current training paradigms where goal-driven data collection creates a dataset bias. In such datasets, language instructions are highly predictable from visual observations alone, causing the conditional mutual information between instructions and actions to vanish, a phenomenon we term Information Collapse. Consequently, models degenerate into vision-only policies that ignore language constraints and fail in out-of-distribution (OOD) settings. To address this, we propose LangForce, a novel framework that enforces instruction following via Bayesian decomposition. By introducing learnable Latent Action Queries, we construct a dual-branch architecture to estimate both a vision-only prior $p(a\mid v)$ and a language-conditioned posterior $\pi(a\mid v,\ell)$. We then optimize the policy to maximize the conditional Pointwise Mutual Information (PMI) between actions and instructions. This objective effectively penalizes the vision shortcut and rewards actions that explicitly explain the language command. Without requiring new data, LangForce significantly improves generalization. Extensive experiments across on SimplerEnv and RoboCasa demonstrate substantial gains, including an 11.3% improvement on the challenging OOD SimplerEnv benchmark, validating the ability of our approach to robustly ground language in action.
LangForce: Bayesian Decomposition of Vision Language Action Models via Latent Action Queries
Shijie Lian1,2,thanks: Equal contribution Bin Yu2,4,11footnotemark: 1 Xiaopeng Lin2,5,11footnotemark: 1 Laurence T. Yang6,1,thanks: Corresponding author Zhaolong Shen2,7 Changti Wu2,8 Yuzhuo Miao2,4 Cong Huang2,3 Kai Chen2,3,9,22footnotemark: 2 1HUST 2ZGCA 3ZGCI 4HIT 5HKUST(GZ) 6ZZU 7BUAA 8ECNU 9DeepCybo
footnotetext: Work done at Zhongguancun Academy (Beijing).
1 Introduction
Vision-Language-Action (VLA) models (Kim et al., 2024; Liu et al., 2025; Bjorck et al., 2025b; Black et al., 2025) have emerged as a promising paradigm for general-purpose robot manipulation, leveraging the vast knowledge of pre-trained Vision-Language Models (VLMs) to ground natural language instructions into physical actions. By training on large-scale datasets of human demonstrations, these models aim to learn a policy $\pi(a\mid v,\ell)$ that can execute diverse tasks specified by language $\ell$ given visual observations $v$.
While demonstrating strong performance in in-distribution settings, current VLA models still face challenges in generalizing to novel instructions or complex multi-task scenarios, particularly in out-of-distribution (OOD) environments (Xing et al., 2025; Xu et al., 2023, 2025). This limitation is especially pronounced during post-training, where fine-tuning on narrow, task-specific datasets can lead to catastrophic forgetting of the VLM’s general capabilities and impair its ability to generalize to new tasks (Xing et al., 2025; Zhou et al., 2025). We hypothesize that this fragility is exacerbated by a prevalent bias in current robotic datasets. Most robotic datasets are collected in a goal-driven manner, where a human operator performs a specific task repeatedly in a fixed scene. In such datasets, the mapping from visual scene $v$ to language instruction $\ell$ is nearly injective: seeing a cabinet in the scene almost invariably implies the task “open the cabinet,” while seeing a bottle implies “pick up the bottle.” This deterministic coupling results in a sharp conditional distribution $p(\ell\mid v)$.
From a Bayesian perspective, the optimal policy can be decomposed as:
$\pi(a\mid v,\ell)=\frac{p(\ell\mid a,v)\,p(a\mid v)}{p(\ell\mid v)}.$ (1)
Here, $p(a\mid v)$ represents a vision-only prior (i.e., what actions are likely in this scene?), and $p(\ell\mid a,v)$ is the likelihood (i.e., how well does action $a$ explain instruction $\ell$?). When $p(\ell\mid v)$ is sharp, the model can predict $\ell$ solely from $v$ without attending to $a$. Consequently, the likelihood term $p(\ell\mid a,v)$ collapses to $p(\ell\mid v)$, and the posterior policy degenerates to the prior:
$\pi(a\mid v,\ell)\approx p(a\mid v).$ (2)
In other words, the model effectively ignores the language instruction, learning a “vision shortcut” that fails whenever the task is ambiguous or the environment changes. We provide a rigorous probabilistic analysis of why this collapse occurs in Appendix A.
To address this, we propose LangForce, a novel framework that explicitly enforces instruction following via Bayesian decomposition. Our key insight is to maximize the conditional Pointwise Mutual Information (PMI) between actions and instructions, which is equivalent to maximizing the log-likelihood ratio (LLR): $\log p(\ell\mid a,v)-\log p(\ell\mid v)$. This objective penalizes the vision shortcut by requiring the action $a$ to provide additional information about $\ell$ that cannot be inferred from $v$ alone.
We instantiate this framework by introducing Latent Action Queries—a set of learnable tokens injected into the VLM. These queries serve a dual purpose: they act as a bottleneck to extract action-relevant features for a downstream Diffusion Transformer (DiT) policy, and they enable a dual-branch training strategy. In the Priori Branch, queries attend only to vision to learn $p(a\mid v)$; in the Posteriori Branch, they attend to both vision and language to learn $\pi(a\mid v,\ell)$. By optimizing the LLR between these branches, LangForce learns to ground language robustly without requiring new data.
Our contributions are three-fold:
  1. 1.
    We identify and empirically validate the “vision shortcut” pathology in current VLA training, showing that standard models often ignore language in favor of dataset-specific visual correlations.
  2. 2.
    We propose LangForce, a method that leverages Latent Action Queries and a dual-branch Bayesian objective to recover the true language-conditioned policy from biased data.
  3. 3.
    We demonstrate that LangForce achieves state-of-the-art performance on SimplerEnv, LIBERO, and RoboCasa, with a remarkable 11.3% improvement in OOD generalization on SimplerEnv, proving its effectiveness in breaking the vision shortcut.
2 Motivation: Vision Shortcut
Before detailing our method, we present empirical evidence to substantiate our hypothesis: that standard VLA models trained on goal-driven datasets often learn a vision-only policy $p(a\mid v)$ rather than a true language-conditioned policy $\pi(a\mid v,\ell)$. Specifically, we employ the Qwen3VL-4B-GR00T model from starVLA (starVLA, 2025) as our representative VLA architecture. We conduct three pilot experiments to reveal this illusion of instruction following. In all three experiments, we train the model by feeding only the visual observation $v$ (masking the language instruction $\ell$), effectively testing the vision-only prior $p(a\mid v)$.
2.1 Experiment 1: The Vision Shortcut in ID Testing
We first train a standard VLA model on a subset of the Humanoid robot tabletop manipulation data from PhysicalAI-Robotics-GR00T-X-Embodiment-Sim (Bjorck et al., 2025b) and evaluate on 24 tasks from the RoboCasa benchmark (Nasiriany et al., 2024). Averaged across all 24 tasks, the vision-only model achieves a success rate of 44.6%, which is close to the language-conditioned baseline of 47.8% (as shown in Table 4). This small gap reveals that the model can succeed without relying on language instructions, as the training and evaluation scenes and tasks are highly similar, enabling the model to learn a near-deterministic mapping from vision to action. Figure 1 provides a relevant example.
Figure 1: Examples of the vision shortcut in RoboCasa (Nasiriany et al., 2024). Training data exhibits visual diversity but limited task diversity. As a result, the model learns to execute tasks directly based on specific visual cues rather than relying on language instructions.
2.2 Experiment 2: Failure in Ambiguous Scenarios
To further investigate this behavior, we train a standard VLA model on the LIBERO benchmark (Liu et al., 2023), which contains four subsets: Spatial, Object, Long, and Goal. We train on all four training sets and evaluate on all four test sets. The vision-only model achieves success rates comparable to the full VLA model on three subsets (Spatial: 90.2%, Object: 99.6%, Long: 86.0% in Vision-Only, Spatial: 97.8%, Object: 98.8%, Long: 92.0% in Baseline), where each visual scene corresponds to a single task. However, on the LIBERO Goal subset, the vision-only success rate plummets to 9.8% (97.4% in Baseline).
Figure 2: (a) In LIBERO Goal (Liu et al., 2023), the same scene corresponds to multiple tasks, revealing ambiguity that vision-only models fail to resolve. (b) Action loss curves on BridgeDataV2 (Walke et al., 2023) and Fractal (Brohan et al., 2022) show the vision-only model achieves comparable loss to the full model, indicating the presence of visual shortcuts even in diverse, in-the-wild datasets.
The key difference is that LIBERO Goal presents inherent ambiguity: multiple valid tasks are associated with the same object configuration during training. For instance, a scene with multiple bowls, a stove, and a drawer could correspond to either “put bowl in drawer” or “put bowl on stove”. This confirms that while the model can exploit vision-action correlations in unambiguous datasets, when multiple tasks share the same visual context, due to a lack of language to resolve ambiguity, the model is dominated by the prior $p(a\mid v)$ learned from dataset statistics. Figure 2(a) illustrates examples where the same visual scene in LIBERO Goal corresponds to multiple distinct tasks.
2.3 Experiment 3: Failure in OOD Generalization
Finally, we test the generalization capability by training on the high-quality BridgeDataV2 dataset (Walke et al., 2023) and Fractal (Brohan et al., 2022) dataset (diverse, in-the-wild scenes) and evaluating on SimplerEnv (Li et al., 2024d) (simulation, OOD). During training on BridgeDataV2, the vision-only model achieves an action loss of 0.13, comparable to the full language-conditioned model’s loss of 0.08 (as shown in Figure 2(b)). This indicates that even in diverse, in-the-wild scenarios, the model can still identify visual shortcuts (e.g., specific lighting or background features mapping to specific actions) to minimize the training objective without truly grounding the language instructions.
However, this reliance on visual shortcuts proves catastrophic for generalization. When evaluated on SimplerEnv, which presents visually distinct simulation environments, the vision-only baseline achieves near 0% success. This confirms that the low training loss on Bridge was achieved by overfitting to domain-specific visual patterns rather than learning generalizable manipulation skills. Consequently, when these specific visual cues are absent in the OOD environment, the policy fails completely.
2.4 Theoretical Insight: Information Collapse
We formalize the “vision shortcut” as a collapse of the conditional mutual information (CMI) between instructions and actions. Ideally, a robust VLA policy should maintain high $I(\ell;a\mid v)$, meaning the action choice significantly reduces uncertainty about the instruction. However, the CMI is bounded by the conditional entropy of the language:
$I(\ell;a\mid v)=H(\ell\mid v)-H(\ell\mid a,v)\leq H(\ell\mid v).$ (3)
In goal-driven datasets, the deterministic mapping $v\to\ell$ implies $H(\ell\mid v)\approx 0$ (Xu et al., 2025). Consequently, $I(\ell;a\mid v)$ is forced to vanish, theoretically preventing the model from learning any dependency between $a$ and $\ell$ beyond what is already captured by $v$. To break this deadlock, we cannot rely on standard imitation learning. Instead, we must explicitly intervene to maximize the information gain provided by the action. This motivates our use of the Log-Likelihood Ratio (LLR), which effectively estimates the Pointwise Mutual Information (PMI), rewarding the policy only when it captures the specific semantics of $\ell$ that are not predictable from $v$.
3 Method: LangForce
In this section, we introduce LangForce, a framework designed to mitigate the vision shortcut in VLA models. We present the overall framework of LangForce in Figure 3. We first formalize the problem through a Bayesian lens (Section 3.1), deriving an objective that maximizes the mutual information between actions and instructions. We then present our architecture, which uses Latent Action Queries to instantiate this decomposition (Section 3.2), and detail our dual-branch training strategy (Section 3.3).
Figure 3: The framework of LangForce. The framework employs a dual-branch architecture with shared VLM weights. The Priori Branch (left) processes $v,\mathcal{Q},\ell]$ with causal masking to learn the vision-only prior $p(a\mid v)$. The Posteriori Branch (right) processes $v,\ell,\mathcal{Q}]$ to learn the full policy $\pi(a\mid v,\ell)$. Latent Action Queries $\mathcal{Q}$ serve as a bottleneck interface, and the LLR objective (in Eq. 7) encourages the model to maximize the information between actions and instructions. At inference, only the Posteriori Branch is used, incurring no additional computational overhead.
3.1 Objective Formulation
As established in Section 2.4, standard VLA training on goal-driven datasets leads to information collapse where $\pi(a\mid v,\ell)\to p(a\mid v)$. To counteract this, we propose to regularize the policy by maximizing the conditional Pointwise Mutual Information (PMI) between the action and the instruction. This objective can be formulated as maximizing the Log-Likelihood Ratio (LLR) between the posterior policy and the vision-only prior:
$\mathcal{L}_{\text{LLR}}=\log\frac{\pi(a\mid v,\ell)}{p(a\mid v)}=\log p(\ell\mid a,v)-\log p(\ell\mid v).$ (4)
The detailed derivation is provided in Appendix B. This formulation requires us to simultaneously model the posterior $\pi(a\mid v,\ell)$ and the prior $p(a\mid v)$. In the following sections, we describe how LangForce efficiently instantiates these two distributions using a shared architecture with Latent Action Queries.
3.2 Latent Action Queries
To efficiently instantiate the proposed Bayesian decomposition within a unified VLM architecture, we introduce Latent Action Queries. We extend the VLM vocabulary with $K=64$ learnable tokens, denoted as $\mathcal{Q}=\{\texttt{<|action\_1|>},\dots,\texttt{<|action\_K|>}\}$. Correspondingly, the VLM’s embedding layer is expanded with $K$ learnable embedding vectors. These vectors are learned to aggregate task-relevant information from the preceding vision and language tokens into a compact latent representation for action execution.
These queries function as a dedicated bottleneck at the interface between the VLM (e.g., Qwen3-VL (Bai et al., 2025)) and the continuous action head (a Diffusion Transformer (Peebles and Xie, 2023)). Crucially, while the VLM typically processes the full sequence, the bottleneck is enforced by exclusively forwarding the hidden states of these query tokens, $\mathbf{H}_{\mathcal{Q}}\in\mathbb{R}^{K\times D}$, to the action expert. This contrasts with recent VLA architectures such as $\pi_{0}$ (Black et al., 2024) and GR00T (Bjorck et al., 2025b; GEAR-Team et al., 2025), which typically feed the hidden states of all input tokens to the action head.
This design choice is critical: by leveraging the causal masking inherent in decoder-only VLMs, we can precisely control the information encoded in $\mathbf{H}_{\mathcal{Q}}$ simply by changing the position of $\mathcal{Q}$ in the input sequence. This flexibility enables the strict separation of vision-only and vision-language contexts required for our dual-branch strategy.
3.3 Dual-Branch Training Framework
We propose a training paradigm with two parallel branches sharing the same VLM weights but different input structures.
1. Priori Branch (Vision-Only).
To estimate the prior $p(a\mid v)$, we construct the input sequence as:
$\text{Input}_{\text{prior}}=[v,\mathcal{Q},\ell].$ (5)
Due to the causal attention mask of the decoder-only VLM, the tokens in $\mathcal{Q}$ can attend to the visual observation $v$ but cannot attend to the language instruction $\ell$ (which appears later). Thus, the hidden states $\mathbf{H}_{\mathcal{Q}}^{\text{prior}}$ encode purely visual information. We use these features to predict the action $a$, optimizing a flow-matching loss $\mathcal{L}_{\text{prior}}$ to learn the dataset’s inherent action bias. Crucially, we detach $\mathbf{H}_{\mathcal{Q}}^{\text{prior}}$ from the computation graph when optimizing $\mathcal{L}_{\text{prior}}$ to ensure that the gradient updates for the vision-only prior are confined to the DiT action head, preventing the shared VLM backbone from learning visual shortcuts.
2. Posteriori Branch (Vision + Language).
To estimate the true policy $\pi(a\mid v,\ell)$, we arrange the input as:
$\text{Input}_{\text{post}}=[v,\ell,\mathcal{Q}].$ (6)
Here, $\mathcal{Q}$ appears after $\ell$, allowing it to attend to both vision and language. The resulting hidden states $\mathbf{H}_{\mathcal{Q}}^{\text{post}}$ encode the full context. We optimize a main flow-matching loss $\mathcal{L}_{\text{main}}$ to learn the expert action.
3. Maximizing the Likelihood Ratio.
In addition to action prediction, we explicitly optimize the LLR objective. We treat the VLM’s language modeling loss as a proxy for $\log p(\ell\mid\dots)$. Specifically, in the Priori Branch, the language tokens $\ell$ attend to $v,\mathcal{Q}]$. Since $\mathcal{Q}$ encodes the action information $a$ (via the prior), the probability of generating $\ell$ in this branch approximates $p(\ell\mid v,a_{\text{prior}})$. In the Posteriori Branch, we can compute a baseline $p(\ell\mid v)$ by detaching gradients. However, a more direct and numerically stable approach is to maximize the difference in log-probabilities of the language tokens between the two branches. We define the LLR loss as:
$\mathcal{L}_{\text{LLR}}=\log p(\ell\mid v,\mathbf{H}_{\mathcal{Q}}^{\text{prior}})-\text{sg}\left(\log p(\ell\mid v)\right),$ (7)
where $\text{sg}(\cdot)$ denotes the stop-gradient operator. We maximize this term (minimize $\mathcal{L}_{\text{LLR}}$) to force the action representations $\mathbf{H}_{\mathcal{Q}}$ to carry information that explains $\ell$. The stop-gradient operation is employed to prevent the model from trivially maximizing the ratio by degrading the baseline $p(\ell\mid v)$ (i.e., damaging the VLM’s general language capabilities) rather than improving the numerator.
3.4 Total Training Objective
We train the action decoder using the Rectified Flow Matching objective (Liu et al., 2022; Bjorck et al., 2025b). Specifically, we apply this objective to both the Priori Branch (conditioned on $\mathbf{H}_{\mathcal{Q}}^{\text{prior}}$) and the Posteriori Branch (conditioned on $\mathbf{H}_{\mathcal{Q}}^{\text{post}}$). Given a condition $\mathbf{C}\in\{\mathbf{H}_{\mathcal{Q}}^{\text{post}},\mathbf{H}_{\mathcal{Q}}^{\text{prior}}\}$, the flow-matching loss is defined as:
[…]
where $v_{\psi}$ is the Diffusion Transformer (DiT) predicting the velocity field, $\mathbf{a}_{1}$ is the ground truth action trajectory, $\mathbf{a}_{0}\sim\mathcal{N}(0,I)$ is sampled from a standard Gaussian, and $\mathbf{a}_{t}=(1-t)\mathbf{a}_{0}+t\mathbf{a}_{1}$ represents the interpolated state at timestep $t\in[0,1]$.
The final training loss combines the action prediction losses from both branches with the LLR maximization term:
$\mathcal{L}_{\text{total}}=(1-\lambda)\mathcal{L}_{\text{FM}}(\psi;\mathbf{H}_{\mathcal{Q}}^{\text{post}})+\lambda\mathcal{L}_{\text{FM}}(\psi;\mathbf{H}_{\mathcal{Q}}^{\text{prior}})-\beta\mathcal{L}_{\text{LLR}},$ (9)
where $\lambda$ balances the contribution of the prior and posterior action losses, and $\beta$ controls the strength of the LLR regularization. We set $\lambda=0.3$ and $\beta=0.1$ in our experiments. During inference, we exclusively execute the Posteriori Branch to obtain $\mathbf{H}_{\mathcal{Q}}^{\text{post}}$ and generate actions via the DiT. This ensures that our method incurs no additional computational overhead compared to standard VLA baselines at test time.
4 Experiment
To comprehensively evaluate the effectiveness of LangForce, we conducted extensive experiments on three simulation benchmarks: SimplerEnv, LIBERO, and RoboCasa. Our training pipeline is built upon the StarVLA framework (starVLA, 2025), distributed across 8 NVIDIA H100 GPUs, and strictly follows its default training protocols to ensure fair comparison. In our experiments, LangForce is instantiated on the QwenGR00T architecture from StarVLA. We employ the AdamW optimizer (Loshchilov and Hutter, 2017) initialized with a learning rate of 1e-5 and a cosine annealing schedule. System-level optimizations include DeepSpeed ZeRO-2 (Rasley et al., 2020), gradient clipping at a norm of 1.0, and no gradient accumulation. All baseline performance metrics are obtained from their original papers or other peer-reviewed publications. To ensure a fair comparison, the training datasets for these baselines encompass the data used in our experiments.
4.1 Experiments on SimplerEnv
We utilize two large-scale subsets from the Open X-Embodiment (OXE) dataset: BridgeDataV2 (Walke et al., 2023) and Fractal (Brohan et al., 2022). The model is fine-tuned for 50k steps on a cluster of 8 GPUs (batch size 16 per device). This benchmark includes four manipulation tasks: “Put spoon on towel”, “Put carrot on plate”, “Stack green cube on yellow cube”, and “Put eggplant in yellow basket”. For each task, we evaluate the VLA policies using the official evaluation scripts provided by the SimplerEnv repository (Li et al., 2024d). To mitigate the effects of randomness, we run 480 independent trials and report the average performance (Avg@480).
Table 1: Results of evaluating the VLA models with the WidowX robot in the SimplerEnv simulation environment. We highlight the best results in bold and the second-best results with underline.
Method Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average RT-1-X (O’Neill et al., 2024) 0.0 4.2 0.0 0.0 1.1 Octo-Base (Team et al., 2024) 15.8 12.5 0.0 41.7 17.5 Octo-Small (Team et al., 2024) 41.7 8.2 0.0 56.7 26.7 OpenVLA-OFT (Kim et al., 2025) 34.2 30.0 30.0 72.5 41.8 RoboVLM (Li et al., 2024c) 50.0 37.5 0.0 83.3 42.7 Magma (Yang et al., 2025a) 37.5 29.2 20.8 91.7 44.8 CogACT (Li et al., 2024b) 71.7 50.8 15.0 67.5 51.3 SpatialVLA (Qu et al., 2025) 20.8 20.8 25.0 70.8 34.4 TraceVLA (Zheng et al., 2025b) 12.5 16.6 16.6 65.0 27.7 VideoVLA (Shen et al., 2025) 75.0 20.8 45.8 70.8 53.1 $\pi_{0}$ (Black et al., 2024) 29.2 62.5 29.2 91.6 53.1 $\pi_{0.5}$ (Black et al., 2025) 49.3 64.7 44.7 69.7 57.1 Isaac-GR00T-N1.6-Bridge (GEAR-Team et al., 2025) 64.5 65.5 5.5 93.0 57.1 QwenGR00T (Baseline) + Qwen3-VL-4B (starVLA, 2025) 87.5 50.0 29.2 54.2 55.2 LangForce + Qwen3-VL-4B 89.6 63.8 33.3 79.2 66.5
The results are summarized in Table 1. LangForce consistently outperforms comparison baselines, achieving a state-of-the-art average success rate of 66.5%. Notably, compared to the direct baseline QwenGR00T (55.2%) built on the same StarVLA framework, our method delivers an absolute improvement of 11.3%, validating that the performance gain stems from our proposed Bayesian decomposition rather than the base architecture. Significant improvements are observed in tasks requiring precise object identification and manipulation, such as “Put Carrot on Plate” (+13.6%) and “Put Eggplant in Yellow Basket” (+15.0%). Furthermore, LangForce surpasses other recent strong competitors, including the flow-matching-based $\pi_{0.5}$ (57.1%) and the dual-system Isaac-GR00T-N1.6 (57.1%). These results confirm that by explicitly optimizing the mutual information between language and action, LangForce effectively mitigates the vision shortcut. Fundamentally, this validates that our approach prevents the policy from collapsing into a spurious vision-only prior $p(a|v)$ caused by dataset determinism, and instead compels the model to learn the true causal dependency of actions on language instructions.
4.2 Experiments on LIBERO
Table 2: Comparison on the LIBERO benchmark. We train one policy for all 4 suites. Avg@500 success rates (%) across four task suites: Spatial, Object, Goal, and Long.
Method Spatial Object Goal Long Avg
OpenVLA (Kim et al., 2024) 87.4 88.4 79.2 53.7 76.5
OpenVLA-OFT (Kim et al., 2025) 97.6 98.4 97.9 94.5 97.1
$\pi_{0}$ (Black et al., 2024) 96.8 98.8 95.8 85.2 94.1
$\pi_{0.5}$ (Black et al., 2025) 98.8 98.2 98.0 92.4 96.9
Qwen3-VL-FAST 97.3 97.4 96.3 90.6 95.4
Qwen3-VL-OFT 97.8 98.6 96.2 93.8 96.6
Qwen3-VL-GR00T 97.8 98.8 97.4 92.0 96.5
VisionOnly Qwen3-VL-GR00T 90.2 99.6 9.8 86.0 71.4
LangForce 99.2 99.6 99.4 95.2 98.4
We also evaluate LangForce on the LIBERO benchmark (Liu et al., 2023). Given that the training and testing environments in LIBERO are highly similar and current VLA research has largely saturated this benchmark (with baselines exceeding 95%), our method yields comparable performance on the Spatial, Object, and Long suites. However, on the Goal suite, LangForce achieves a success rate of 99.4%, outperforming the Qwen3-VL-GR00T baseline (97.4%) by +2.0%. As highlighted in Section 2.2, the Goal suite features significant visual ambiguity where multiple tasks share the same scene. This result empirically validates that our method effectively mitigates the vision shortcut, enabling the model to resolve ambiguity through robust instruction following.
To further substantiate this claim, we conducted a quantitative analysis of the conditional entropy $H(\ell\mid v)$ on the LIBERO Goal dataset. We approximated this metric by computing the Negative Log-Likelihood (NLL) of the ground-truth instructions given the visual observations across 40,000 samples. The results are summarized in Table 3.
Table 3: Quantitative analysis of conditional entropy on LIBERO Goal. We report the Negative Log-Likelihood (NLL) and Perplexity (PPL) of instructions given visual observations, serving as a proxy for $H(\ell\mid v)$. Higher NLL/PPL indicates that the model preserves the necessary uncertainty about the task given only vision, preventing the information collapse observed in baselines.
Method NLL (nats/token) $\uparrow$ PPL ($\exp(\text{NLL})$) $\uparrow$ Std Dev (NLL)
QwenGR00T (Baseline) 8.51 4964.1 0.55
LangForce 9.47 12964.9 0.54
LangForce achieves a significantly higher NLL (9.47 nats/token) compared to the QwenGR00T baseline (8.51 nats/token). This difference is even more pronounced in terms of Perplexity (PPL), where LangForce reaches 12964.9 compared to the baseline’s 4964.1. It is worth noting that the language structure in LIBERO Goal is highly repetitive (e.g., “put the [object] in/on the [receptacle]”), which naturally encourages the model to fit these syntactic patterns. Despite this, our method achieves a higher NLL and PPL, indicating that the uncertainty stems primarily from the task-specific nouns (objects and receptacles) rather than the sentence structure. In the context of LIBERO Goal, where visual scenes are inherently ambiguous, the baseline’s lower NLL implies it is “overconfident” in predicting these key nouns solely from vision. This confirms that standard training leads to a collapse in conditional entropy, where the model learns spurious correlations ($v\to\ell$) and ignores the actual need for language disambiguation. In contrast, the higher NLL observed in LangForce indicates that our model prevents the pathological collapse of $H(\ell\mid v)$ observed in standard training, preserving uncertainty levels that align with the inherent ambiguity of visual scenes. By maintaining this necessary uncertainty, our method forces the policy to actively utilize the provided language instruction to resolve ambiguity, directly contributing to the superior performance on the Goal suite.
4.3 Experiments on RoboCasa
We evaluate our method on the RoboCasa GR1 Tabletop Manipulation Benchmark (Nasiriany et al., 2024), which consists of 24 diverse manipulation tasks. These tasks feature complex interactions with articulated objects and varied geometries, exemplified by specific tasks like “PnPBottleToCabinetClose” and “PnPCanToDrawerClose”, as well as scenarios involving appliances like microwaves and toasters. For training, we utilize the Humanoid Robot Tabletop Manipulation subset from the PhysicalAI-Robotics-GR00T-X-Embodiment-Sim (Bjorck et al., 2025b) dataset. All other settings follow Section 4.1. To guarantee statistical significance, we evaluate each task using 50 independent trials and report the average success rate (Avg@50).
Table 4: Results of evaluating the VLA models with the GR1 robot in the RoboCasa Tabletop simulation environment. The results for Isaac-GR00T N1.5 and Isaac-GR00T N1.6 are sourced from the official Isaac-GR00T github repository (Bjorck et al., 2025b). The results for QwenGR00T, QwenPI, QwenOFT, QwenFAST are sourced from the official starVLA experiments (starVLA, 2025). We highlight the best results in bold and the second-best results with underline.
Task Isaac-GR00T N1.5 Isaac-GR00T N1.6 QwenGR00T +Qwen3VL QwenPI +Qwen3VL QwenOFT +Qwen3VL QwenFAST +Qwen3VL VisionOnly QwenGR00T LangForce +Qwen3VL PnP Bottle To Cabinet Close 54.0 51.5 46.0 26.0 30.0 38.0 70.0 72.0 PnP Can To Drawer Close 50.0 13.0 80.0 62.0 76.0 44.0 78.0 78.0 PnP Cup To Drawer Close 38.0 8.5 54.0 42.0 44.0 56.0 42.0 46.0 PnP Milk To Microwave Close 60.0 14.0 48.0 50.0 44.0 44.0 50.0 56.0 PnP Potato To Microwave Close 32.0 41.5 28.0 42.0 32.0 14.0 44.0 36.0 PnP Wine To Cabinet Close 38.0 16.5 46.0 32.0 36.0 14.0 40.0 46.0 PnP * to * Close (Avg) 45.3 24.2 50.3 42.3 43.7 35.0 54.0 55.7 PnP Novel From Cuttingboard To Basket 38.0 58.0 48.0 40.0 50.0 54.0 58.0 66.0 PnP Novel From Cuttingboard To Cardboardbox 46.0 46.5 40.0 46.0 40.0 42.0 26.0 40.0 PnP Novel From Cuttingboard To Pan 58.0 68.5 68.0 60.0 70.0 58.0 72.0 68.0 PnP Novel From Cuttingboard To Pot 62.0 65.0 52.0 40.0 54.0 58.0 50.0 48.0 PnP Novel From Cuttingboard To Tieredbasket 28.0 46.5 56.0 44.0 38.0 40.0 20.0 44.0 PnP Novel From Cuttingboard To * (Avg) 46.4 56.9 52.8 46.0 50.4 50.4 45.2 53.2 PnP Novel From Placemat To Basket 30.0 58.5 42.0 44.0 32.0 36.0 48.0 54.0 PnP Novel From Placemat To Bowl 60.0 57.5 44.0 52.0 58.0 38.0 32.0 62.0 PnP Novel From Placemat To Plate 56.0 63.0 48.0 50.0 52.0 42.0 34.0 52.0 PnP Novel From Placemat To Tieredshelf 36.0 28.5 18.0 28.0 24.0 18.0 16.0 24.0 PnP Novel From Placemat To * (Avg) 45.5 51.9 38.0 43.5 41.5 33.5 32.5 48.0 PnP Novel From Tray To Cardboardbox 52.0 51.5 38.0 34.0 44.0 28.0 50.0 50.0 PnP Novel From Tray To Plate 48.0 71.0 56.0 64.0 56.0 34.0 64.0 58.0 PnP Novel From Tray To Pot 60.0 64.5 50.0 44.0 62.0 46.0 52.0 62.0 PnP Novel From Tray To Tieredbasket 52.0 57.0 36.0 50.0 54.0 36.0 42.0 44.0 PnP Novel From Tray To Tieredshelf 32.0 31.5 16.0 28.0 30.0 16.0 16.0 22.0 PnP Novel From Tray To * (Avg) 48.8 55.1 39.2 44.0 49.2 32.0 44.8 47.2 PnP Novel From Plate To Bowl 58.0 57.0 60.0 52.0 60.0 52.0 26.0 54.0 PnP Novel From Plate To Cardboardbox 44.0 43.5 50.0 40.0 50.0 30.0 38.0 48.0 PnP Novel From Plate To Pan 60.0 51.0 54.0 36.0 66.0 48.0 44.0 54.0 PnP Novel From Plate To Plate 64.0 78.7 70.0 48.0 68.0 50.0 60.0 78.0 PnP Novel From Plate To * (Avg) 56.5 57.6 58.5 44.0 61.0 45.0 42.0 58.5 Average 48.2 47.6 47.8 43.9 48.8 39.0 44.7 52.6
The quantitative results on RoboCasa are presented in Table 4. LangForce achieves a state-of-the-art average success rate of 52.6%, surpassing all competing baselines including QwenOFT (48.8%), Isaac-GR00T N1.5 (48.2%), and the direct baseline QwenGR00T (47.8%). Notably, our method substantially outperforms the VisionOnly baseline (44.7%), demonstrating the effectiveness of language grounding. Consistent with the empirical evidence presented in our motivation (Section 2.1), the VisionOnly baseline achieves a surprisingly high success rate of 44.7%, lagging only slightly behind the standard QwenGR00T baseline (47.8%). This observation reconfirms the prevalence of the vision shortcut in this benchmark, suggesting that a significant portion of tasks can be solved by relying solely on visual cues. However, LangForce breaks this performance ceiling, achieving a state-of-the-art average success rate of 52.6% and surpassing all competing baselines, including QwenOFT (48.8%), Isaac-GR00T N1.5 (48.2%), and Isaac-GR00T N1.6 (47.6%). Crucially, our method demonstrates substantial gains in tasks where the vision-only policy falters. For instance, in “PnP Novel From Placemat To Bowl”, LangForce reaches 62.0%, substantially surpassing both the VisionOnly baseline (32.0%) and QwenGR00T (44.0%). These results indicate that maximizing the LLR objective successfully forces the policy to extract and utilize task-specifying information from language, rather than settling for local optima based on visual shortcuts.
Figure 4: Qualitative comparison of general multimodal reasoning. We present a case where the model is asked to solve a mathematical problem. The standard VLA baseline (QwenGR00T) suffers from catastrophic forgetting; while the text before the comma implies “differentiating all terms together”, the subsequent output degenerates into repetitive and meaningless gibberish (bottom right). In contrast, LangForce (top right) retains the VLM’s original reasoning and language generation capabilities (left), successfully solving the problem.
Figure 5: Additional qualitative comparison. Demonstrating the preservation of general VLM capabilities on another example.
4.4 Preservation of General Capabilities
A prevalent concern in the field is that fine-tuning VLMs for robotic action generation (VLA training) often compromises the model’s foundational reasoning and multimodal understanding, leading to a degradation of general conversational abilities (Zhou et al., 2025; Xu et al., 2025; Hancock et al., 2025; Yu et al., 2026). ChatVLA (Zhou et al., 2025) attributes this to spurious forgetting, where robot training overwrites critical visual-text alignments, and task interference between control and understanding objectives.
We observe similar degradation in our baseline: as illustrated in Figure 4 and Figure 5, the standard QwenGR00T model loses the ability to converse coherently, even when prompted with pure text inputs. In stark contrast, LangForce remarkably preserves these capabilities when queried with language instructions. This is because our LLR objective enforces a strong dependency on language for action generation, preventing the linguistic representations from collapsing or becoming redundant.
It is important to qualify these findings with a nuanced distinction: while LangForce effectively preserves the model’s text-only conversational abilities, its general vision-language capabilities (involving both image and text inputs) may still experience some degradation following VLA training. We attribute this to our training setup, which requires the vision tower and multimodal projection layers to adapt for specific control tasks, inevitably shifting visual representations away from the original pre-trained alignment manifold. Such specialization appears to be a natural trade-off when optimizing an embodied agent for precise action execution.
Crucially, however, LangForce successfully prevents the catastrophic collapse of the backbone’s text-only language behavior. In standard baselines, the prevalence of vision shortcuts often renders instruction tokens effectively redundant for control. This redundancy weakens the gradient signal for language processing, leading to a drift in shared parameters that manifests as failures even on pure text queries (as seen in Fig. 4 and Fig. 5). In contrast, our LLR objective acts as a powerful regularizer by enforcing a strong, explicit dependency on language. This mechanism maintains the functional utility of instruction tokens, thereby preserving the backbone’s text-only conversational ability even as the visual modality specializes for control. This retention of the linguistic core is of significant practical value: it ensures that the VLM does not degenerate into a shallow feature mapper, but rather preserves its potential for higher-level reasoning and generalization—key motivations for employing Foundation Models in robotics.
4.5 Ablation Studies
We conduct ablation studies on SimplerEnv to validate the contributions of individual components in LangForce. All experiments utilize the Qwen3-VL-4B backbone, and results are presented in Table 5.
Effectiveness of Bayesian Decomposition.
Comparing the full LangForce (63.5%) with the “+ Action Query” ablation (57.5%), we observe a significant performance boost (+6.0%). This indicates that while the architectural changes provide some benefit, the core improvement stems from our dual-branch Bayesian learning objective. By explicitly modeling and maximizing the pointwise mutual information (PMI) between instructions and actions, the model effectively overcomes the vision shortcut, validating the central hypothesis of this work. Additionally, we provide ablation studies on the hyperparameters $\lambda$ and $\beta$ in Appendix C.1.
Table 5: Ablation study on SimplerEnv. All experiments are based on the Qwen3-VL-4B backbone. We compare the baseline QwenGR00T, the addition of Latent Action Queries, and the full LangForce framework to validate the contributions of each component.
Method Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average QwenGR00T (starVLA, 2025) 87.5 50.0 29.2 54.2 55.2 QwenGR00T + Action Query 74.6 58.3 29.2 67.9 57.5 LangForce 89.6 63.8 33.3 79.2 66.5
Potential of Latent Action Queries.
Even without the dual-branch definition, introducing Latent Action Queries (“+ Action Query”) improves upon the QwenGR00T baseline (55.2% $\to$ 57.5%). This suggests that Latent Action Queries function as a promising architectural inductive bias. Unlike standard approaches that feed full sequences of vision and language token embeddings into the action decoder, our query-based mechanism forces the VLM to compress and summarize task-relevant information into a compact set of latent tokens. From a computational perspective, this design is highly efficient. It decouples the complexity of the Diffusion Transformer (DiT) from the length of the VLM input context. Specifically, the complexity of condition processing in the DiT is reduced from $O(N^{2})$ (scaling with the massive number of vision-language tokens $N$) to $O(K^{2})$ (scaling with the small, constant number of query tokens $K$), thereby streamlining the action generation process. We also investigate the impact of the number of latent action queries on model performance. Detailed analysis and results are provided in Appendix C.2.
5 Discussion
Based on our analysis of the vision shortcut and the Bayesian decomposition framework, we discuss several potential insights that may guide future research and community practices.
Rethinking Data Collection Strategies.
Our experiments suggest that the deterministic mapping from visual scenes to language instructions ($H(\ell\mid v)\approx 0$) in goal-driven datasets is a significant factor contributing to the vision shortcut. To mitigate this, we hypothesize that a shift in data collection strategies could be beneficial. Prioritizing data collection in ambiguous scenarios—where the task cannot be inferred solely from the initial observation—might naturally increase the conditional entropy of language. By enriching datasets with scenes that support multiple valid tasks, models may be forced to rely more heavily on instructions for disambiguation. Consequently, we advocate for:
Prioritizing ambiguous scenarios during data collection to naturally compel models to rely on language for disambiguation.
Leveraging Human Data for Robustness.
Recently, there has been growing interest in training robot models on large-scale human video data, such as HRDT (Bi et al., 2025), In-N-On (Cai et al., 2025), METIS (Fu et al., 2025), and PhysBrain (Lin et al., 2025). Unlike curated robot datasets, human activities are inherently multimodal and context-dependent; the same environment often hosts a wide variety of behaviors, potentially leading to a less sharp $p(\ell\mid v)$. This observation highlights a critical opportunity for scaling up robot learning:
Leveraging human video data, with its rich behavioral diversity, to mitigate the information collapse typical of curated robot datasets.
World Models as an Alternative Bayesian Formulation.
Beyond the VLM framework focused on in this work, recent studies have also explored adapting World Models for VLA control, as seen in F1-VLA (Lv et al., 2025), Mantis (Yang et al., 2025b), and InternVLA-A1 (contributors, 2026). From a theoretical perspective, these approaches can be viewed as an alternative instantiation of the Bayesian rule, specifically performing inverse dynamics on imagined futures. If we consider $v$ as a sequence of past frames $v_{\leq t}$, and treat the future state $v_{t+1}$ as a latent variable generated by the model (conditioned on $\ell$), the action inference can be expressed as:
$p(a\mid v_{\leq t},v_{t+1},\ell)=\frac{p(v_{t+1}\mid v_{\leq t},a,\ell)\,p(a\mid v_{\leq t},\ell)}{p(v_{t+1}\mid v_{\leq t},\ell)}.$ (10)
Here, the numerator $p(v_{t+1}\mid v_{\leq t},a,\ell)$ represents a world model (forward dynamics) predicting the future state. The term $p(a\mid v_{\leq t},\ell)$ serves as an action prior, and the denominator $p(v_{t+1}\mid v_{\leq t},\ell)$ represents the future prediction marginalized over actions. In this formulation, the policy execution involves first “imagining” a desired future $v_{t+1}$ consistent with $\ell$, and then inferring the optimal action $a$ via the equation above. Since world models are typically trained on vast amounts of video data, the predictive distribution (the numerator) is often rich and highly sensitive to the action $a$. We hypothesize that this sensitivity prevents the collapse of the numerator to the denominator, offering a robust technical path for future research:
Exploring World Models as an alternative Bayesian formulation, where the sensitivity of forward dynamics inherently resists the vision shortcut.
6 Related Work
We build our work upon the following rigorous foundations:
Vision-Language-Action Dataset and Benchmark. The advancement of generalist robot policies relies heavily on large-scale datasets and rigorous benchmarks. LIBERO (Liu et al., 2023) pioneered the systematic study of knowledge transfer in lifelong robot learning. To scale up real-world data, BridgeData V2 (Walke et al., 2023) provided diverse interaction trajectories on low-cost hardware. This effort was expanded by Open X-Embodiment (OXE) (O’Neill et al., 2024), which aggregated data across 22 robot embodiments, and Droid (Khazatsky et al., 2024), which further increased diversity with distributed data collection. For scalable evaluation, RoboCasa (Nasiriany et al., 2024) introduced a large-scale simulation framework with realistic kitchen environments, while SimplerEnv (Li et al., 2024d) provided a simulated evaluation proxy to correlate with real-world performance, addressing the reproducibility crisis in physical evaluation. More recently, RoboTwin 2.0 (Chen et al., 2025) offered a unified benchmark for bimanual manipulation with automated data generation, and AgiBot-World (Bu et al., 2025) scaled training data to over 1 million trajectories with human-in-the-loop verification.
Vision-Language-Action Models. To bridge the gap between semantic understanding and physical control, Vision-Language-Action (VLA) models have emerged as a dominant paradigm. Early works like Octo (Team et al., 2024) established the transformer-based policy as a versatile initialization, utilizing diffusion heads to handle multimodal action distributions. Building on pre-trained VLMs, OpenVLA (Kim et al., 2024) and its optimized variant OpenVLA-OFT (Kim et al., 2025) fine-tune large language models for robotic control, demonstrating strong generalization.
To further enhance action modeling, specialized architectures have been proposed. CogACT (Li et al., 2024b) and RDT-1B (Liu et al., 2025) employ dedicated diffusion transformers to capture complex action dynamics. Recent works explore advanced architectures and generation algorithms to better balance reasoning and control. The GR00T series (N1, N1.5, N1.6) (Bjorck et al., 2025b, a; GEAR-Team et al., 2025) utilizes a dual-system design, architecturally coupling a VLM for perception with a diffusion head for action generation. In parallel, the $\pi_{0}$ series ($\pi_{0}$, $\pi_{0.5}$, FAST) (Black et al., 2024, 2025; Pertsch et al., 2025) leverages Flow Matching as a superior generation objective, adopting a two-stage training paradigm that pre-trains the VLM backbone before fine-tuning. Specifically, FAST introduces discrete motion tokens to efficiently encode actions for the VLM, while $\pi_{0.5}$ incorporates subtask decomposition to enhance long-horizon planning.
Other approaches like X-VLA (Zheng et al., 2025a) introduce embodiment-specific soft prompts to facilitate cross-embodiment generalization. By learning separate sets of embeddings for each data source, X-VLA effectively leverages heterogeneous robot data with minimal additional parameters. SpatialVLA (Qu et al., 2025) argues that spatial understanding is central to manipulation, introducing Ego3D Position Encoding and Adaptive Action Grids to inject 3D information and learn transferable spatial action knowledge. Then, VideoVLA (Shen et al., 2025) explores transforming video generation models into robot manipulators. By jointly predicting action sequences and future visual outcomes, it leverages the “visual imagination” of generative models to enhance generalization across novel tasks and objects. TwinBrainVLA (Yu et al., 2026) mitigates catastrophic forgetting by coordinating a frozen generalist “Left Brain” for semantic understanding and a trainable specialist “Right Brain” for sensorimotor learning, effectively balancing high-level reasoning with low-level control. BayesVLA (Xu et al., 2025) also employs Bayesian decomposition to improve instruction following. It utilizes a two-stage framework: first training a vision-conditioned prior on large-scale vision-action pairs, and then freezing this prior to train a language-conditioned posterior. However, this design strictly requires a separated training process where the vision component must be frozen in the second stage. In contrast, LangForce enables single-stage, end-to-end training. Furthermore, BayesVLA has not open-sourced its code, which regrettably prevents a direct empirical comparison. Overall, other VLA approaches lack systematic solutions to the vision shortcut problem, where models ignore language instructions in goal-driven datasets.
7 Conclusion
In this work, we identified a critical pathology in standard VLA training: the vision shortcut, where models learn to ignore language instructions in goal-driven datasets, leading to a collapse of the mutual information between instructions and actions. This phenomenon results in catastrophic failure when the model faces ambiguous scenarios or out-of-distribution environments where visual shortcuts no longer hold. To address this, we introduced LangForce, a novel framework that formalizes VLA learning through a Bayesian lens. By decomposing the policy into a vision-only prior and a language-conditioned posterior, we derived a training objective that explicitly maximizes the Log-Likelihood Ratio (LLR), effectively optimizing the Pointwise Mutual Information (PMI) between language and action. We instantiated this decomposition using a unified architecture with Latent Action Queries and a dual-branch training strategy, which incurs no additional computational overhead during inference. Our extensive experiments on SimplerEnv, LIBERO and RoboCasa demonstrate that LangForce effectively mitigates the vision shortcut problem. It significantly outperforms baselines under various conditions; in out-of-distribution simulation environments, where standard methods fail completely, LangForce achieves robust generalization. These results highlight the importance of introducing structural inductive biases in learning robust, generalizable robot policies. We hope our work inspires further research into the information-theoretic foundations of embodied AI, pushing towards robots that truly understand and reliably execute human intent.
Limitation and Future Work
While LangForce offers significant improvements in robustness, the dual-branch architecture introduces a limitation regarding computational overhead during training. Since the model must compute both the Priori and Posteriori branches, the computational cost per iteration theoretically increases. However, we note that the visual input prefix is identical for both branches, and the number of visual tokens vastly outnumbers that of the language and latent action query tokens. By employing a prefix prefill strategy to compute and reuse the visual representations (e.g., vision encoder outputs) for both branches, the actual increase in training time is marginal. Thus, the additional computational overhead remains within a completely acceptable range.
For future work, we intend to provide a more comprehensive empirical evaluation in subsequent versions of this manuscript. We plan to extend our experimental validation to include the RoboTwin2 (Chen et al., 2025) and BEHAVIOR-1K (Li et al., 2024a) benchmarks, alongside real-world robot experiments. We also aim to scale our framework to larger foundation models, such as conducting experiments with LangForce on Qwen3VL-8B or Qwen3VL-30B-A3B (Bai et al., 2025).
Appendix A Probabilistic Analysis of the Vision Shortcut
We analyze the theoretical conditions under which the Bayesian ratio term becomes uninformative, leading the policy to collapse into a vision-only shortcut. Recall the ratio term defined in Eq. 1:
$\mathcal{R}(a,\ell,v)\triangleq\frac{p(\ell\mid a,v)}{p(\ell\mid v)}.$ (11)
Here, the Ratio is Conditional PMI. By applying Bayes’ rule, the log-ratio can be rewritten as:
$\log\mathcal{R}(a,\ell,v)=\log p(\ell\mid a,v)-\log p(\ell\mid v)=\log\frac{p(a,\ell\mid v)}{p(a\mid v)\,p(\ell\mid v)}.$ (12)
This is precisely the Pointwise Mutual Information (PMI) between the action $a$ and language $\ell$, conditioned on the visual observation $v$. Taking the expectation over the data distribution $p(v,\ell,a)$ yields the Conditional Mutual Information (CMI):
$\mathbb{E}_{p(v,\ell,a)}\left[\log\mathcal{R}(a,\ell,v)\right]=I(\ell;a\mid v).$ (13)
This insight reveals that the Bayesian ratio $\mathcal{R}$ provides a meaningful learning signal if and only if the action $a$ carries information about the instruction $\ell$ that is not already contained in $v$. In other words, optimization on $\mathcal{R}$ requires non-negligible $I(\ell;a\mid v)$.
Information Collapse under Deterministic Labeling.
In many goal-driven robotic datasets, the environment setup $v$ often strongly implies the task $\ell$ (e.g., a dirty pot implies cleaning). Mathematically, this means the conditional entropy $H(\ell\mid v)$ approaches zero. Using the entropy identity for CMI:
$I(\ell;a\mid v)=H(\ell\mid v)-H(\ell\mid a,v)\leq H(\ell\mid v).$ (14)
Consequently, as $H(\ell\mid v)\to 0$, the upper bound on the mutual information vanishes, forcing $I(\ell;a\mid v)\to 0$ (Xu et al., 2025). In this regime, $\mathbb{E}[\log\mathcal{R}]\approx 0$, rendering the specific LLR objective uninformative on average for samples drawn from such biased distributions. This theoretical hypothesis is strongly supported by our empirical results in Section 2.1, where a Vision-Only policy achieves a high success rate of 44.7% on RoboCasa, nearly matching the instruction-following baseline (47.8%). This confirms that for a significant portion of the data, $v$ indeed suffices to infer the task.
Language becomes Unnecessary for Imitation Learning.
When the dataset exhibits the property $\ell\approx f(v)$ (i.e., single-instruction annotation per scene), the demonstration distribution satisfies:
$p_{\text{data}}(a\mid v,\ell)\approx p_{\text{data}}(a\mid v).$ (15)
Under standard Behavior Cloning (minimizing $\mathcal{L}_{\text{BC}}=\mathbb{E}_{p_{\text{data}}}[-\log\pi_{\theta}(a\mid v,\ell)]$), an optimal policy $\pi^{*}$ exists that depends solely on vision: $\pi^{*}(a\mid v,\ell)=p_{\text{data}}(a\mid v)$. Crucially, in this regime, the training objective provides no gradient incentive to capture the semantic dependency on $\ell$, as the vision-only solution already minimizes the loss globally. This creates a “vision shortcut”, where the model learns to ignore instructions. While reliable in-distribution, this policy fails catastrophically in OOD settings where visual ambiguity increases ($H(\ell\mid v)$ becomes large) and language becomes necessary to disambiguate the task.
Appendix B Derivation of the LLR Objective
Building upon the probabilistic analysis in Appendix A, we now derive the practical learning objective used in LangForce to counteract the vision shortcut. Our core motivation is to maximize the learning signal discussed previously: the Conditional Pointwise Mutual Information (PMI) between the action $a$ and the language instruction $\ell$, given the visual observation $v$. The PMI is formally defined as:
$\text{PMI}(a,\ell\mid v)=\log\frac{\pi(a,\ell\mid v)}{p(a\mid v)p(\ell\mid v)}.$ (16)
By applying the chain rule of probability $\pi(a,\ell\mid v)=\pi(a\mid v,\ell)p(\ell\mid v)$, we can rewrite the PMI as the log-ratio between the posterior policy and the vision-only prior:
$\text{PMI}(a,\ell\mid v)=\log\frac{\pi(a\mid v,\ell)p(\ell\mid v)}{p(a\mid v)p(\ell\mid v)}=\log\frac{\pi(a\mid v,\ell)}{p(a\mid v)}.$ (17)
This formulation highlights that maximizing PMI is equivalent to maximizing the divergence between the language-conditioned policy and the vision-only prior, directly penalizing the collapse where $\pi(a\mid v,\ell)\approx p(a\mid v)$.
Alternatively, using the chain rule $\pi(a,\ell\mid v)=p(\ell\mid a,v)p(a\mid v)$, we arrive at the second form, which constitutes our practical LLR objective:
$\text{PMI}(a,\ell\mid v)=\log\frac{p(\ell\mid a,v)p(a\mid v)}{p(a\mid v)p(\ell\mid v)}=\log\frac{p(\ell\mid a,v)}{p(\ell\mid v)}=\log p(\ell\mid a,v)-\log p(\ell\mid v).$ (18)
This objective represents the difference between the log-likelihood of the instruction given the action and vision, and the log-likelihood of the instruction given vision alone. Maximizing this quantity compels the model to select actions $a$ that make the instruction $\ell$ significantly more probable than determining it from the visual context $v$ alone, thereby extracting the additional information required to solve ambiguous or out-of-distribution tasks.
Appendix C Additional Experiments
C.1 Ablation on $\lambda$ and $\beta$ in Loss Function
Table 6: Ablation study on $\beta$ in Eq. 9. All experiments are based on the Qwen3-VL-4B backbone.
$\lambda$ Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average 0 86.4 64.6 26.0 76.0 63.3 0.1 84.2 65.6 27.3 81.9 64.8 0.3 89.6 63.8 33.3 79.2 66.5 0.5 81.2 61.5 25.8 87.7 64.1
Table 7: Ablation study on $\beta$ in Eq. 9. All experiments are based on the Qwen3-VL-4B backbone.
$\beta$ Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average 0 80.8 61.0 30.8 72.5 61.3 0.1 89.6 63.7 33.3 79.2 66.5 0.2 82.5 67.5 34.2 77.9 65.5 0.3 73.8 58.3 29.2 85.2 61.6
We analyze the impact of hyperparameters $\lambda$ and $\beta$ based on the results in Table 6 and Table 7. First, regarding the prior loss weight $\lambda$, we observe that even when $\lambda=0$ (where the prior branch is used solely for LLR calculation without explicit action supervision), the model achieves an average success rate of 63.3%. This represents a substantial improvement over the QwenGR00T baseline (55.2%), confirming that the LLR objective alone effectively regularizes the policy. Increasing $\lambda$ to 0.3 yields the optimal performance of 66.5%, validating that explicitly learning the vision-only prior $p(a|v)$ further aids the decomposition. The performance remains robust across $\lambda\in[0,0.5]$, though it drops slightly at higher values.
Second, for the LLR weight $\beta$, setting $\beta=0$ corresponds to training the dual-branch architecture without the mutual information maximization term. This configuration achieves 61.3%, surpassing the baseline (55.2%). This suggests that the architectural design, which explicitly separates a vision-only pathway to absorb dataset biases, inherently helps the posterior branch focus on language. Incorporating the LLR term ($\beta=0.1$) further boosts the success rate to 66.5%.
C.2 Ablation on Number of Latent Action Queries
To explore the potential of Latent Action Queries as a out-of-the-box technique, we conduct this ablation solely on the query mechanism, without incorporating our proposed Bayesian decomposition. We analyse the impact of the number of latent action queries on model performance, as summarized in Table 8. We observe that increasing the number of queries from 16 to 64 leads to a substantial improvement in success rate. However, doubling the queries to 128 results in performance saturation without providing additional benefits. Consequently, we adopt 64 queries as the optimal configuration to balance performance and computational overhead.
Table 8: Ablation study on Number of Latent Action Query. All experiments are based on the Qwen3-VL-4B backbone.
Number Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average 16 69.8 47.9 20.8 60.4 49.7 32 62.5 57.4 25.8 79.2 56.2 64 74.6 58.3 29.2 67.9 57.5 128 72.9 57.3 29.1 70.8 57.5